EVA Report – April 3rd

Crew 208 EVA Report 03-04-2019

EVA # 5

Author: David Kim

Purpose of EVA: Geologic exploration

Start time: 1030h

End time: 1300h

Narrative: Drove Curiosity and Spirit to destination through planned route. Arrived at Gray Moon and walked around area collecting Martian soil samples and took pictures. Returned without issues. Noted that Spirit originally drained its battery quickly from 98% to 80% in less than 5 minutes during normal operation. However, this normalized and drained slower after 80%. All normal rules and operations followed (Back Wheel Drive, Normal speed, Normal gear).

Destination: x516500, y4254500

Coordinates: x516500, y4254500

Participants: Amanda Manget (ENG), David Kim (HSO), Erika Rydberg (GHO)

Road(s) and routes per MDRS Map: Cow Dung Road 0110, to Rd #1572, to Rd #1575

Mode of travel: Spirit and Curiosity

Operations Report – April 3rd

Crew 208 Operations Report 03-APR-2019

SOL: 6

Name of person filing report: Amanda Manget

Non-nominal systems: Generator

Notes on non-nominal systems: Radiator hoses and alternator need replacement on the generator. A contractor has been arranged to make these repairs within a few days.

Generator (hours run): 16

Solar – SOC 65% @ 1730 (before David turned on generator for the night)

Diesel Reading – 45%

Propane Reading – 51%

Ethanol Free Gasoline – 2 gallons

Water (auxiliary tank) – 0 gallons (Note: Not in Use)

Water (static tank) – 179.79

Auxiliary to Static tank transfer – No

Gallons transferred: 0

Water in GreenHab – 203 gallons

Static to Loft Pump used – Yes

Water Meter: 0142405.8 gallons @ 1700

Toilet tank emptied: Yes

Deimos rover used: No, not functional

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Spirit rover used: Yes

Hours: 86

Beginning charge: 98%

Ending charge: 45%

Currently charging: Yes

Opportunity rover used: No

Hours: 55.7

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Curiosity rover used: Yes

Hours: 86.9

Beginning charge: 100%

Ending charge: 42%

Currently charging: Yes

Notes on rovers: Spirit started at 98% even though it had been charged overnight. 5 minutes into our EVA, Spirit dropped down to 80%, then decreased at a normal rate thereafter. When it reached 65%, we returned to the Hab.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3)

Reason for use:

Oil Added? No

ATV Fuel Used: 0 Gals

# Hours the ATVs were Used today: 0

Notes on ATVs: None

HabCar used and why, where? Not Used

CrewCar used and why, where? Not Used

CrewCar odometer reading: 175109 miles

General notes and comments: Nothing to report

Summary of internet: Nothing to report

Summary of suits and radios: EV suits normal, all batteries read > 12.5 volts

Summary of Hab operations: Checked SOC of electrical system as follows:

0730 SOC 100%, generator turned off by David

1000 SOC 98%, generator off

1230 SOC 90%, generator off

1730 SOC 65%, generator turned on by David

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAMM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: None

Greenhab Report – April 3rd

GreenHab Officer: Erika Rydberg

Environmental control:

Ambient, door open.

Shade cloth (80%) on

Average temperatures:

Average: 25.5 C

Low: 20 C
High: 31.4 C

Hours of supplemental light: 0

Daily water usage for crops: 5 gallons

Water in Blue Tank— 203 gallons

Time(s) of watering for crops: 8:45 and 17:30

Changes to crops: Continued to remove new basil flowers to keep basil growing.

Narrative: It warmed up in the Greenhab today, kept a close eye on the temperature.

Harvest: None

Support/supplies needed: None

Astronomy Report – April 3rd

Astronomy Report

Name: Amanda Manget

Crew: 208
Date: 03-APR-2019


Robotic Telescope Requested: MDRS-WF

Objects to be Imaged this Evening (if weather permits):

Submitted March 27th: Pinwheel (active 7/10 images taken)

Submitted April 1st: Crab Nebula, Dumbbell Nebula

Objects Already Imaged: Pinwheel with MDRS-14, Whirlpool Galaxy with MDRS-WF

Problems Encountered: High winds last night causing dome to stay closed. If weather permits, images will be taken tonight.


Solar Features Observed: N/A

Problems Encountered: N/A

Journalist Report – April 3rd

[title Journalist Report – April 3rd]

Crew 208 Journalist Report

Date: April 3rd 2019

Author: Erika Rydberg

Making It Work On Mars

Today we had a number of successes in our experience and travels here on Mars. We successfully printed another 3D map, we were able to fix our non-functioning 3D printer, and Dean and myself had solid one-hour rides on our bicycle to charge the battery we’re using to power our 3D printer.

David, Amanda, and I had an awesome EVA out to check out an area north of the Hab, where the landscape is called Grey Moon. Did it look moon-like? Well we haven’t visited there yet ourselves, but it was surprising and the color description was accurate. We hoped to find fossils, but instead found only the breathtaking landscape in front of us, which was enough to satisfy us on our journey. The longer we stay, the more the landscape continues to be surprising, surreal, and varied.

A short report for today, but as usual we’re never quite sure what tomorrow will hold.

Sol Summary – April 3rd

Crew 208 Sol Summary Report 03-APR-2019

Sol: 6

Summary Title: Geology Rocks! (in sim)

Author’s name: Julielynn Wong

Mission Status: Nominal

Sol Activity Summary:

Crew 208 continued 3D printing in the Hab and worked on their research manuscripts.

Useful things to 3D print at MDRS:

Ice cube tray

ENG + HSO + GHO used the Curiosity and Spirit rovers during their EVA to search for fossils.

Crew XO + HSO tried out our Space VR goggles.

Crew XO + GHO each exercised for 1 hour to generate power for our bike-powered 3D printing study.

Look Ahead Plan: Continue 3D printing in Hab and data collection for thoracentesis study in Science Dome. Plan to break sim on Sol 8.

Anomalies in work: None

Weather: Sunny

Crew Physical Status: Nominal

EVA: #5 to Grey Moon and Beige Moon

Reports to be filed: Operations, Greenhab, Journalist, EVA Report, EVA Request

Support Requested: None

Copyright © The Mars Society. All rights reserved. | Main Site